# Transport, Variational Inference and Diffusions: with Applications to Annealed Flows and Schrödinger Bridges

### v1

@@ E.4 Drift based EM
$=-\tfrac{\sigma^{2}}{2}\int_{0}^{T}|\nabla\phi_{t}|^{2}({\bm{Y}}_{t})\,\mathrm{d}t+\sigma\int_{0}^{T}\nabla\phi_{t}({\bm{Y}}_{t})\cdot\mathrm{d}{\bm{W}}_{t}=\operatorname{ln}\left(\frac{\mathrm{d}{\mathbb{P}}^{\phi}}{\mathrm{d}{\mathbb{P}}}\right),$ (E.5)
@@
The claim now follows, since the unique solution to the Schrödinger problem is characterised by the product-form expression in (66, see Léonard [2014a, Section 2], together with the marginal constraints ${\mathbb{P}}^{\phi}_{0}=\mu$ and ${\mathbb{P}}^{\phi}_{T}=\nu$, which are satisfied by assumption. ∎
Remark 15 (Connection to reciprocal classes [Shi et al., 2023, Peluchetti, 2023] and TD learning [Liu et al., b]).
The calculation in equation (E.5) makes the relationship between the HJB equation (28) and reciprocal classes manifest (since reciprocal classes can essentially be defined through the relationship (66), see Léonard et al. [2014], Rœlly [2013]). Moreover, equation (E.5) showcases the relationship between TD learning [Sutton and Barto, 2018, Chapter 6] as suggested in Liu et al. [b] and HJB regularisation. Indeed,
@@
where the variance is taken with respect to the path measure induced by (63), is a valid HJB-regulariser in the sense of Corollary 4.4. The equivalence between $\mathrm{Reg}_{\mathrm{BSDE}}(\phi)=0$ and the HJB equation (28) follows from the theory of backward stochastic differential equations (BSDEs)121212… not to be confused with reverse-time SDEs as in (2.2)., see, for example, the proof of Proposition 3.4 in Nüsken and Richter [2023] and the discussion in Nüsken and Richter [2021, Section 3.2].
In the following, we present an analogue of Proposition 4.3 involving the backward drift [Chen et al., 2019]:
@@
As opposed to Chen et al. [2016, equation (41)], the HJB-equation (69) does not involve the time reversal of the Schrödinger prior; the form of the HJB equations is not uniquely determined. On the other hand, (69) contains the divergence term $\nabla\cdot f$, which discourages us from enforcing this constraint in the same way as (28). An akin result can be found in [Liu et al., b] stated in terms of BSDEs.
@@
Using the forward-backward Radon-Nikodym derivative in (2.2), we compute
@@
$-\sigma\int_{0}^{T}\theta_{t}({\bm{Y}}_{t})\cdot\overleftarrow{\mathrm{d}}{\bm{W}}_{t}+\tfrac{\sigma^{2}}{2}\int_{0}^{T}|\nabla\theta_{t}|^{2}({\bm{Y}}_{t})\,\mathrm{d}t$
$=\operatorname{ln}\left(\frac{\mathrm{d}\mu}{\mathrm{d}\mathrm{Leb}}\right)-\operatorname{ln}\left(\frac{\mathrm{d}\nu}{\mathrm{d}\mathrm{Leb}}\right)-\sigma\int_{0}^{T}(\nabla\cdot f_{t})({\bm{Y}}_{t})\,\mathrm{d}t-\sigma\int_{0}^{T}\theta_{t}({\bm{Y}}_{t})\cdot\overleftarrow{\mathrm{d}}{\bm{W}}_{t}+\tfrac{\sigma^{2}}{2}\int_{0}^{T}|\nabla\theta_{t}|^{2}({\bm{Y}}_{t})\,\mathrm{d}t.$

### v6

@@ E.5 HJB-Regularizers
$=-\tfrac{\sigma^{2}}{2}\int_{0}^{T}|\nabla\phi_{t}|^{2}({\bm{Y}}_{t})\,\mathrm{d}t+\sigma\int_{0}^{T}\nabla\phi_{t}({\bm{Y}}_{t})\cdot\mathrm{d}{\bm{W}}_{t}=\operatorname{ln}\left(\frac{\mathrm{d}{\mathbb{P}}^{\phi}}{\mathrm{d}{\mathbb{P}}}\right),$ (79)
@@
The claim now follows, since the unique solution to the Schrödinger problem is characterised by the product-form expression in (80, see Léonard (2014a, Section 2), together with the marginal constraints ${\mathbb{P}}^{\phi}_{0}=\mu$ and ${\mathbb{P}}^{\phi}_{T}=\nu$, which are satisfied by assumption. ∎
Remark 12 (Summarised relationship to previous work).
For $\lambda=0$, coordinate-wise updates of $\mathcal{L}_{\mathrm{Schr}}(\phi,\theta)$ recover the IPF updates from De Bortoli et al. (2021); Vargas et al. (2021a) according to Corollary 3.1. Note that $\mathcal{L}_{\mathrm{Schr}}$ is an unconstrained objective, in contrast to (17); previous works (Koshizuka & Sato, 2023; Zhang & Katsoulakis, 2023) have suggested incorporating the marginal constraints softly by adding penalising terms to the running cost in (17). Those approaches require a limiting argument (from an algorithmic standpoint, adaptive tuning of a weight parameter) to recover the solution to (17). In contrast, the conclusion of Proposition E.3 holds for arbitrary $\lambda>0$. Shi et al. (2023); Peluchetti (2023) suggest an algorithm involving reciprocal projections onto the reciprocal class associated to $f_{t}$. From Clark (1991); Thieullen (2002); Rœlly (2013), the HJB-equation (76) is a local characteristic ($\mathrm{Reg}(\phi)=0$ forces (75) to be in the reciprocal class); hence $\mathrm{Reg}(\phi)$ in (74) plays a similar role as the reciprocal projection (Shi et al., 2023, Definition 3), see Remark 13. Liu et al. (a) suggest an iterative IPF-like scheme involving a temporal difference term (Sutton & Barto, 2018, Chapter 6). As in Nüsken & Richter (2023), this is a an HJB-regulariser in the sense of Proposition E.3, see Remark 13. Finally, Albergo et al. (2023, Theorem 5.3) and Gushchin et al. (2022) develop saddle-point objectives for (17).
Remark 13 (Connection to reciprocal classes (Shi et al., 2023; Peluchetti, 2023) and TD learning (Liu et al., a)).
The calculation in equation (79) makes the relationship between the HJB equation (76) and reciprocal classes manifest (since reciprocal classes can essentially be defined through the relationship (80), see Léonard et al. (2014); Rœlly (2013)). Moreover, equation (79) showcases the relationship between TD learning (Sutton & Barto, 2018, Chapter 6) as suggested in Liu et al. (a) and HJB regularisation. Indeed,
@@
where the variance is taken with respect to the path measure induced by (77), is a valid HJB-regulariser in the sense of Proposition E.3. The equivalence between $\mathrm{Reg}_{\mathrm{BSDE}}(\phi)=0$ and the HJB equation (76) follows from the theory of backward stochastic differential equations (BSDEs)121212… not to be confused with reverse-time SDEs as in (12)., see, for example, the proof of Proposition 3.4 in Nüsken & Richter (2023) and the discussion in Nüsken & Richter (2021, Section 3.2).
In the following, we present an analogue of Proposition E.4 involving the backward drift (Chen et al., 2019):
@@
As opposed to Chen et al. (2016, equation (41)), the HJB-equation (83) does not involve the time reversal of the Schrödinger prior; the form of the HJB equations is not uniquely determined. On the other hand, (83) contains the divergence term $\nabla\cdot f$, which discourages us from enforcing this constraint in the same way as (76). An akin result can be found in Liu et al. (a) stated in terms of BSDEs.
@@
Using the forward-backward Radon-Nikodym derivative in (14), we compute
@@
$-\sigma\int_{0}^{T}\nabla\theta_{t}({\bm{Y}}_{t})\cdot\overleftarrow{\mathrm{d}}{\bm{W}}_{t}+\tfrac{\sigma^{2}}{2}\int_{0}^{T}|\nabla\theta_{t}|^{2}({\bm{Y}}_{t})\,\mathrm{d}t$
$=\operatorname{ln}\left(\frac{\mathrm{d}\mu}{\mathrm{d}\mathrm{Leb}}\right)\!-\!\operatorname{ln}\left(\frac{\mathrm{d}\nu}{\mathrm{d}\mathrm{Leb}}\right)-\sigma\!\!\int_{0}^{T}\!\!\!\!\!(\nabla\cdot f_{t})({\bm{Y}}_{t})\,\mathrm{d}t-\sigma\int_{0}^{T}\!\!\!\!\!\!\nabla\theta_{t}({\bm{Y}}_{t})\cdot\overleftarrow{\mathrm{d}}{\bm{W}}_{t}+\tfrac{\sigma^{2}}{2}\!\!\int_{0}^{T}\!\!\!|\nabla\theta_{t}|^{2}({\bm{Y}}_{t})\,\mathrm{d}t.$